Ben Blum recently graduated from Moonshadow Universitys accounting program. He has been hired as an analyst by Primrose TireElimination of North East Lost sales revenue Less: avoidable variable expenses Lost contribution margin
Answer #1

Elimination of north east

Lost sales revenue

(200000)

Less: avoidable variable expenses (120000+50000)

170000

Lost contribution margin

$(30000)

Variable cost of goods sold = 153000-33000 = 120000

Variable operating costs = 96000-46000 = 50000